# Model-assisted inference for dynamic causal effects in staggered rollout cluster randomized experiments

### v2

@@ 4 Comparisons and recommendations
We compare different regression estimators regarding their efficiency in estimating $\tau_{j}(a,a^{\prime})$ and $\theta$ and provide relevant recommendations. For ease of exposition, we first focus on the individual DWATE estimand before returning to the summary estimands in Section 4.4. The theoretical results of regression estimators, marginal standard errors, and corresponding theorems are summarized in Table 3.
@@ 4.1 Estimators using individual-level data and cluster-period averages
We first provide additional results on the connections between estimators using individual-level data and cluster-period averages and summarize them in Proposition 1. Here, we explicitly write specific types of covariates in brackets after an estimator to emphasize the covariates used for baseline adjustment. All covariates are centered when included in the regression model. For example, $\widehat{\tau}_{j,\mathrm{I}}^{\mathrm{adj}}(a,a^{\prime})[\bm{X}_{ijk}]$ denotes the regression estimator using individual-level data adjusted for $\bm{X}_{ijk}$ as covariates in the working model.
@@ 4.2 Estimators using individual-level data and scaled cluster-period totals
The inequality $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})[\pi_{ij\cdot},\widetilde{\bm{X}}_{ij\cdot}]\}\leq\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})[\pi_{ij\cdot}]\}$ is established following the techniques to prove the result that covariate adjustment will always improve the asymptotic variance under the independent data setting (Lin, 2013). The inequality $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})[\pi_{ij\cdot}]\}\leq\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{I}}(a,a^{\prime})\}$ can be obtained by first establishing the following expression $\sum_{i=1}^{I}\widetilde{\epsilon}_{ij\cdot}(a)^{2}=\sum_{i=1}^{I}\{\widetilde{Y}_{ij\cdot}(a)-I\pi_{ij\cdot}\overline{Y}_{\cdot j\cdot}(a)\}^{2}=\sum_{i=1}^{I}\{\widetilde{Y}_{ij\cdot}(a)-\overline{Y}_{\cdot j\cdot}(a)-(\pi_{ij\cdot}-I^{-1})I\overline{Y}_{\cdot j\cdot}(a)\}^{2}$. Then, by the Frisch-Waugh-Lovell Theorem (Frisch and Waugh, 1933), the quantity $\sum_{i=1}^{I}r_{ij}(a)^{2}$ for $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})[\pi_{ij\cdot}]\}$ in Theorem 6 is the residual sum of squares from the OLS fit of $\widetilde{Y}_{ij\cdot}(a)-\overline{Y}_{ij\cdot}(a)$ on $\pi_{ij\cdot}-I^{-1}$, which by definition is smaller than $\sum_{i=1}^{I}\{\widetilde{Y}_{ij\cdot}(a)-\overline{Y}_{\cdot j\cdot}(a)-(\pi_{ij\cdot}-I^{-1})I\overline{Y}_{\cdot j\cdot}(a)\}^{2}$. Therefore, adjusting for $\pi_{ij\cdot}$ in the regression using scaled cluster-period totals improves the simple difference-in-means, and adjusting for scaled cluster-period totals of additional covariates leads to greater efficiency improvement asymptotically. The same argument can be applied to obtaining the inequality $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})[\pi_{ij\cdot},\widetilde{\bm{X}}_{ij\cdot}]\}\leq\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{I}}^{\mathrm{adj}}(a,a^{\prime})\}$. Specifically, the quantity $\sum_{i=1}^{I}r_{ij}(a)^{2}=\sum_{i=1}^{I}\{\widetilde{\epsilon}_{ij\cdot}(a)-\widetilde{\bm{X}}_{ij\cdot}^{c}\bm{\gamma}_{j,\mathrm{I}}(a)\}^{2}$ for $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{I}}^{\mathrm{adj}}(a,a^{\prime})\}$ in Theorem 2 can be expressed as $\sum_{i=1}^{I}\{\widetilde{Y}_{ij\cdot}(a)-\overline{Y}_{\cdot j\cdot}(a)-(\pi_{ij\cdot}-I^{-1})I\overline{Y}_{\cdot j\cdot}(a)-\widetilde{\bm{X}}_{ij\cdot}^{c}\bm{\gamma}_{j,\mathrm{I}}(a)\}^{2}$, which is greater than the residual sum of squares from the OLS fit of $\widetilde{Y}_{ij\cdot}(a)-\overline{Y}_{ij\cdot}(a)$ on $(\pi_{ij\cdot}-I^{-1},\widetilde{\bm{X}}_{ij\cdot}^{c})$, that is, the quantity $\sum_{i=1}^{I}r_{ij}(a)^{2}$ for $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})[\pi_{ij\cdot},\widetilde{\bm{X}}_{ij\cdot}]\}$ in Theorem 6. Therefore, adjusting for $(\pi_{ij\cdot},\widetilde{\bm{X}}_{ij\cdot})$ in the regression using scaled cluster-period totals improves the regression adjustment of $\bm{X}_{ijk}$ using the individual-level data, regarding asymptotic efficiency.
@@ 4.3 Estimators using cluster-period averages and scaled cluster-period totals
Proposition 3 is obtained via similar arguments centered around the sum of squares and the Frisch-Waugh-Lovell Theorem used in deriving Proposition 2. The inequality $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})[\pi_{ij\cdot},\pi_{ij\cdot}\bm{C}_{ij}]\}\leq\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})[\pi_{ij\cdot}]\}$ follows the result that regression adjustment will always improve the asymptotic variance (Lin, 2013). The inequality $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})[\pi_{ij\cdot}]\}\allowbreak\leq\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{A}}(a,a^{\prime})\}$ is obtained from the fact that $\sum_{i=1}^{I}\widetilde{\epsilon}_{ij\cdot}(a)^{2}=\sum_{i=1}^{I}\{\widetilde{Y}_{ij\cdot}(a)-I\pi_{ij\cdot}\overline{Y}_{\cdot j\cdot}(a)\}^{2}=\sum_{i=1}^{I}\{\widetilde{Y}_{ij\cdot}(a)-\overline{Y}_{\cdot j\cdot}(a)-(\pi_{ij\cdot}-I^{-1})I\overline{Y}_{\cdot j\cdot}(a)\}^{2}$, and $\sum_{i=1}^{I}r_{ij}(a)^{2}$ for $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})[\pi_{ij\cdot}]\}$ in Theorem 6 is the residual sum of squares from the OLS fit of $\widetilde{Y}_{ij\cdot}(a)-\overline{Y}_{ij\cdot}(a)$ on $\pi_{ij\cdot}-I^{-1}$, which by definition is smaller than $\sum_{i=1}^{I}\{\widetilde{Y}_{ij\cdot}(a)-\overline{Y}_{\cdot j\cdot}(a)-(\pi_{ij\cdot}-I^{-1})I\overline{Y}_{\cdot j\cdot}(a)\}^{2}$. Therefore, adjusting for $\pi_{ij\cdot}$ improves the efficiency of simple difference-in-means, and adjusting for scaled cluster-period totals of additional covariates leads to greater efficiency improvement. The same argument can be applied to obtain the inequality $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})[\pi_{ij\cdot},\pi_{ij\cdot}\bm{C}_{ij}]\}\leq\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{A}}^{\mathrm{adj}}(a,a^{\prime})\}$, where the quantity $\sum_{i=1}^{I}r_{ij}(a)^{2}=\sum_{i=1}^{I}\{\widetilde{\epsilon}_{ij\cdot}(a)-I\pi_{ij\cdot}\bm{C}_{ij}^{c}\bm{\gamma}_{j,\mathrm{I}}(a)\}^{2}$ for $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{A}}^{\mathrm{adj}}(a,a^{\prime})\}$ in Theorem 4 can be expressed as $\sum_{i=1}^{I}\{\widetilde{Y}_{ij\cdot}(a)-\overline{Y}_{\cdot j\cdot}(a)-(\pi_{ij\cdot}-I^{-1})I\overline{Y}_{\cdot j\cdot}(a)-\pi_{ij\cdot}\bm{C}_{ij}^{c}I\bm{\gamma}_{j,\mathrm{A}}(a)\}^{2}$, which is greater than the residual sum of squares from the OLS fit of $\widetilde{Y}_{ij\cdot}(a)-\overline{Y}_{ij\cdot}(a)$ on $(\pi_{ij\cdot}-I^{-1},\pi_{ij\cdot}\bm{C}_{ij}^{c})$, i.e., the quantity $\sum_{i=1}^{I}r_{ij}(a)^{2}$ for $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})[\pi_{ij\cdot},\pi_{ij\cdot}\bm{C}_{ij}]\}$ in Theorem 6. Therefore, adjusting for $(\pi_{ij\cdot},\pi_{ij\cdot}\bm{C}_{ij})$ improves the regression adjustment of $\bm{C}_{ij}$ using cluster-period averages, regarding asymptotic efficiency.
@@ 4.4 Generalization to summary estimands
Results in Propositions 1 - 3 can be generalized to summary estimands. The generalization of Proposition 1 is apparent. For Propositions 2 and 3, consider the example of comparing $\mathrm{se}^{2}(\widehat{\theta}_{\mathrm{T}}^{\mathrm{adj}}[\pi_{ij\cdot}])$ and $\mathrm{se}^{2}(\widehat{\theta}_{\mathrm{A}}^{\mathrm{adj}})$, where $\mathrm{se}^{2}(\widehat{\theta}_{\mathrm{A}}^{\mathrm{adj}})-\mathrm{se}^{2}(\widehat{\theta}_{\mathrm{T}}^{\mathrm{adj}}[\pi_{ij\cdot}])=\bm{b}_{\theta}^{T}\{\mathbf{Cov}(\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{adj}})-\mathbf{Cov}(\widehat{\bm{\tau}}_{\mathrm{T}}^{\mathrm{adj}}[\pi_{ij\cdot}])\}\bm{b}_{\theta}$. Using the arguments for Proposition 3, the diagonal entries of $\mathbf{Cov}(\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{adj}})$ are greater than the corresponding entries of $\mathbf{Cov}(\widehat{\bm{\tau}}_{\mathrm{T}}^{\mathrm{adj}}[\pi_{ij\cdot}])$. Since both $\mathbf{Cov}(\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{adj}})$ and $\mathbf{Cov}(\widehat{\bm{\tau}}_{\mathrm{T}}^{\mathrm{adj}}[\pi_{ij\cdot}])$ are covariance matrices and, therefore, positive definite, the difference matrix $\mathbf{Cov}(\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{adj}})-\mathbf{Cov}(\widehat{\bm{\tau}}_{\mathrm{T}}^{\mathrm{adj}}[\pi_{ij\cdot}])$ is positive semi-definite such that $\bm{b}_{\theta}^{T}\{\mathbf{Cov}(\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{adj}})-\mathbf{Cov}(\widehat{\bm{\tau}}_{\mathrm{T}}^{\mathrm{adj}}[\pi_{ij\cdot}])\}\bm{b}_{\theta}\geq 0$ for any $\bm{b}_{\theta}\in\mathbb{R}^{J^{2}(J+1)/2}$, suggesting that $\widehat{\theta}_{\mathrm{T}}^{\mathrm{adj}}[\pi_{ij\cdot}]$ is more efficient than $\widehat{\theta}_{\mathrm{A}}^{\mathrm{adj}}$ with $\mathrm{se}^{2}(\widehat{\theta}_{\mathrm{T}}^{\mathrm{adj}}[\pi_{ij\cdot}])\leq\mathrm{se}^{2}(\widehat{\theta}_{\mathrm{A}}^{\mathrm{adj}})$. The following Corollary 2 summarizes the results comparing asymptotic variances of estimators for summary estimands.

### v3

@@ 4 Comparisons and recommendations
We compare different regression estimators in terms of their efficiency in estimating $\tau_{j}(a,a^{\prime})$ and $\theta$, and provide relevant recommendations. For ease of exposition, we first focus on the individual DWATE estimand before returning to the summary estimands in Section 4.4. The theoretical results of regression estimators, marginal standard errors, and corresponding theorems are summarized in Table 3.
@@ 4.1 Estimators using individual-level data and cluster-period averages
We first present additional results on the connections between estimators using individual-level data and cluster-period averages, summarizing them in Proposition 1. Here, we explicitly write specific types of covariates in brackets after an estimator to emphasize the covariates used for baseline adjustment. All covariates are centered when included in the regression model. For example, $\widehat{\tau}_{j,\mathrm{I}}^{\mathrm{adj}}(a,a^{\prime})[\bm{X}_{ijk}]$ denotes the regression estimator using individual-level data adjusted for $\bm{X}_{ijk}$ as covariates in the working model.
@@ 4.2 Estimators using individual-level data and scaled cluster-period totals
The inequality $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})[\pi_{ij\cdot},\widetilde{\bm{X}}_{ij\cdot}]\}\leq\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})[\pi_{ij\cdot}]\}$ is established following the techniques to prove the result that covariate adjustment will always improve the asymptotic variance under the independent data setting (Lin, 2013). The inequality $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})[\pi_{ij\cdot}]\}\leq\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{I}}(a,a^{\prime})\}$ can be obtained by first establishing the following expression $\sum_{i=1}^{I}\widetilde{\epsilon}_{ij\cdot}(a)^{2}=\sum_{i=1}^{I}\{\widetilde{Y}_{ij\cdot}(a)-I\pi_{ij\cdot}\overline{Y}_{\cdot j\cdot}(a)\}^{2}=\sum_{i=1}^{I}\{\widetilde{Y}_{ij\cdot}(a)-\overline{Y}_{\cdot j\cdot}(a)-(\pi_{ij\cdot}-I^{-1})I\overline{Y}_{\cdot j\cdot}(a)\}^{2}$. Then, by the Frisch-Waugh-Lovell Theorem (Frisch and Waugh, 1933), the quantity $\sum_{i=1}^{I}r_{ij}(a)^{2}$ for $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})[\pi_{ij\cdot}]\}$ in Theorem 6 is the residual sum of squares from the OLS fit of $\widetilde{Y}_{ij\cdot}(a)-\overline{Y}_{ij\cdot}(a)$ on $\pi_{ij\cdot}-I^{-1}$, which by definition is smaller than $\sum_{i=1}^{I}\{\widetilde{Y}_{ij\cdot}(a)-\overline{Y}_{\cdot j\cdot}(a)-(\pi_{ij\cdot}-I^{-1})I\overline{Y}_{\cdot j\cdot}(a)\}^{2}$. Therefore, adjusting for $\pi_{ij\cdot}$ in the regression using scaled cluster-period totals improves the simple difference-in-means. Additionally, adjusting for scaled cluster-period totals of other covariates leads to greater asymptotic efficiency improvement. The same argument can be applied to obtaining the inequality $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})[\pi_{ij\cdot},\widetilde{\bm{X}}_{ij\cdot}]\}\leq\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{I}}^{\mathrm{adj}}(a,a^{\prime})\}$. Specifically, the quantity $\sum_{i=1}^{I}r_{ij}(a)^{2}=\sum_{i=1}^{I}\{\widetilde{\epsilon}_{ij\cdot}(a)-\widetilde{\bm{X}}_{ij\cdot}^{c}\bm{\gamma}_{j,\mathrm{I}}(a)\}^{2}$ for $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{I}}^{\mathrm{adj}}(a,a^{\prime})\}$ in Theorem 2 can be expressed as $\sum_{i=1}^{I}\{\widetilde{Y}_{ij\cdot}(a)-\overline{Y}_{\cdot j\cdot}(a)-(\pi_{ij\cdot}-I^{-1})I\overline{Y}_{\cdot j\cdot}(a)-\widetilde{\bm{X}}_{ij\cdot}^{c}\bm{\gamma}_{j,\mathrm{I}}(a)\}^{2}$, which is greater than the residual sum of squares from the OLS fit of $\widetilde{Y}_{ij\cdot}(a)-\overline{Y}_{ij\cdot}(a)$ on $(\pi_{ij\cdot}-I^{-1},\widetilde{\bm{X}}_{ij\cdot}^{c})$, that is, the quantity $\sum_{i=1}^{I}r_{ij}(a)^{2}$ for $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})[\pi_{ij\cdot},\widetilde{\bm{X}}_{ij\cdot}]\}$ in Theorem 6. Therefore, adjusting for $(\pi_{ij\cdot},\widetilde{\bm{X}}_{ij\cdot})$ in the regression using scaled cluster-period totals improves the regression adjustment of $\bm{X}_{ijk}$ using the individual-level data, regarding asymptotic efficiency.
@@ 4.3 Estimators using cluster-period averages and scaled cluster-period totals
Proposition 3 is obtained via similar arguments centered around the sum of squares and the Frisch-Waugh-Lovell Theorem used in deriving Proposition 2. The inequality $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})[\pi_{ij\cdot},\pi_{ij\cdot}\bm{C}_{ij}]\}\leq\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})[\pi_{ij\cdot}]\}$ follows the result that regression adjustment will always improve the asymptotic variance (Lin, 2013). The inequality $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})[\pi_{ij\cdot}]\}\allowbreak\leq\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{A}}(a,a^{\prime})\}$ is obtained from the fact that $\sum_{i=1}^{I}\widetilde{\epsilon}_{ij\cdot}(a)^{2}=\sum_{i=1}^{I}\{\widetilde{Y}_{ij\cdot}(a)-I\pi_{ij\cdot}\overline{Y}_{\cdot j\cdot}(a)\}^{2}=\sum_{i=1}^{I}\{\widetilde{Y}_{ij\cdot}(a)-\overline{Y}_{\cdot j\cdot}(a)-(\pi_{ij\cdot}-I^{-1})I\overline{Y}_{\cdot j\cdot}(a)\}^{2}$, and $\sum_{i=1}^{I}r_{ij}(a)^{2}$ for $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})[\pi_{ij\cdot}]\}$ in Theorem 6 is the residual sum of squares from the OLS fit of $\widetilde{Y}_{ij\cdot}(a)-\overline{Y}_{ij\cdot}(a)$ on $\pi_{ij\cdot}-I^{-1}$, which by definition is smaller than $\sum_{i=1}^{I}\{\widetilde{Y}_{ij\cdot}(a)-\overline{Y}_{\cdot j\cdot}(a)-(\pi_{ij\cdot}-I^{-1})I\overline{Y}_{\cdot j\cdot}(a)\}^{2}$. Therefore, adjusting for $\pi_{ij\cdot}$ improves the efficiency of simple difference-in-means, and adjusting for scaled cluster-period totals of additional covariates leads to a greater improvement in efficiency. The same argument can be applied to obtain the inequality $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})[\pi_{ij\cdot},\pi_{ij\cdot}\bm{C}_{ij}]\}\leq\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{A}}^{\mathrm{adj}}(a,a^{\prime})\}$, where the quantity $\sum_{i=1}^{I}r_{ij}(a)^{2}=\sum_{i=1}^{I}\{\widetilde{\epsilon}_{ij\cdot}(a)-I\pi_{ij\cdot}\bm{C}_{ij}^{c}\bm{\gamma}_{j,\mathrm{I}}(a)\}^{2}$ for $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{A}}^{\mathrm{adj}}(a,a^{\prime})\}$ in Theorem 4 can be expressed as $\sum_{i=1}^{I}\{\widetilde{Y}_{ij\cdot}(a)-\overline{Y}_{\cdot j\cdot}(a)-(\pi_{ij\cdot}-I^{-1})I\overline{Y}_{\cdot j\cdot}(a)-\pi_{ij\cdot}\bm{C}_{ij}^{c}I\bm{\gamma}_{j,\mathrm{A}}(a)\}^{2}$, which is greater than the residual sum of squares from the OLS fit of $\widetilde{Y}_{ij\cdot}(a)-\overline{Y}_{ij\cdot}(a)$ on $(\pi_{ij\cdot}-I^{-1},\pi_{ij\cdot}\bm{C}_{ij}^{c})$, i.e., the quantity $\sum_{i=1}^{I}r_{ij}(a)^{2}$ for $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})[\pi_{ij\cdot},\pi_{ij\cdot}\bm{C}_{ij}]\}$ in Theorem 6. Therefore, adjusting for $(\pi_{ij\cdot},\pi_{ij\cdot}\bm{C}_{ij})$ improves the regression adjustment of $\bm{C}_{ij}$ using cluster-period averages, regarding asymptotic efficiency.
@@ 4.4 Generalization to summary estimands
Results in Propositions 1 - 3 can be generalized to summary estimands. The generalization of Proposition 1 is apparent. For Propositions 2 and 3, consider the example of comparing $\mathrm{se}^{2}(\widehat{\theta}_{\mathrm{T}}^{\mathrm{adj}}[\pi_{ij\cdot}])$ and $\mathrm{se}^{2}(\widehat{\theta}_{\mathrm{A}}^{\mathrm{adj}})$, where $\mathrm{se}^{2}(\widehat{\theta}_{\mathrm{A}}^{\mathrm{adj}})-\mathrm{se}^{2}(\widehat{\theta}_{\mathrm{T}}^{\mathrm{adj}}[\pi_{ij\cdot}])=\bm{b}_{\theta}^{\top}\{\mathbf{Cov}(\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{adj}})-\mathbf{Cov}(\widehat{\bm{\tau}}_{\mathrm{T}}^{\mathrm{adj}}[\pi_{ij\cdot}])\}\bm{b}_{\theta}$. Using the arguments for Proposition 3, the diagonal entries of $\mathbf{Cov}(\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{adj}})$ are greater than the corresponding entries of $\mathbf{Cov}(\widehat{\bm{\tau}}_{\mathrm{T}}^{\mathrm{adj}}[\pi_{ij\cdot}])$. Since both $\mathbf{Cov}(\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{adj}})$ and $\mathbf{Cov}(\widehat{\bm{\tau}}_{\mathrm{T}}^{\mathrm{adj}}[\pi_{ij\cdot}])$ are covariance matrices and, therefore, positive definite, the difference matrix $\mathbf{Cov}(\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{adj}})-\mathbf{Cov}(\widehat{\bm{\tau}}_{\mathrm{T}}^{\mathrm{adj}}[\pi_{ij\cdot}])$ is positive semi-definite such that $\bm{b}_{\theta}^{\top}\{\mathbf{Cov}(\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{adj}})-\mathbf{Cov}(\widehat{\bm{\tau}}_{\mathrm{T}}^{\mathrm{adj}}[\pi_{ij\cdot}])\}\bm{b}_{\theta}\geq 0$ for any $\bm{b}_{\theta}\in\mathbb{R}^{J^{2}(J+1)/2}$, suggesting that $\widehat{\theta}_{\mathrm{T}}^{\mathrm{adj}}[\pi_{ij\cdot}]$ is more efficient than $\widehat{\theta}_{\mathrm{A}}^{\mathrm{adj}}$ with $\mathrm{se}^{2}(\widehat{\theta}_{\mathrm{T}}^{\mathrm{adj}}[\pi_{ij\cdot}])\leq\mathrm{se}^{2}(\widehat{\theta}_{\mathrm{A}}^{\mathrm{adj}})$. The following Corollary 2 summarizes the results comparing asymptotic variances of estimators for summary estimands.